The length of a rectangle is twice the difference of the width and 2. The perimeter is 100 centimeters. What are the width and the length of the rectangle.
Please solve for me
Thank you so much if you solve this I really need help

Answer by addingup(3677)

Twice the difference of the width and 2:
Length is 2 (W-2)
The perimeter is 100:
2L+2W= 100. Since L= 2 (W-2):
2 (2 (W-2))+2W=100
4 (W-2)+2W=100open
4W-8+2W=100 add W on left and add 8 on both sides:
6W= 108 divide both sides by 6:
W=108/6
W= 18 And the length:
2 (18-2)= 32
Check:
2 (18)+2 (32)=100
36+64=100
100=100 we have the correct answer